# Stochastic Deep Learning Surrogate Models for Uncertainty Propagation in Microstructure-Properties of Ceramic Aerogels

### v1

@@ 5 Conclusions
This study introduces an integrated deep learning surrogate modeling framework for efficient uncertainty propagation across the synthesis, microstructure, and properties of ceramic aerogel porous materials. The stochastic microstructure surrogate employs a scalable Wasserstein Generative Adversarial Network with Gradient Penalty (WGAN-GP) to generate ensembles of morphologically consistent microstructures over arbitrarily large domains, trained on smaller domain size samples from Lattice Boltzmann simulations of the foaming process. For property prediction, a Bayesian Convolutional Neural Network (BayesCNN) is developed to estimate weight parameter distributions using small, uncertain high-fidelity datasets from finite element elasticity simulations. Numerical results demonstrate that the WGAN-GP effectively generates microstructures consistent with the morphology of training data (in terms of pore size distribution and two-point correlation) across larger domains. The BayesCNN accurately maps microstructural images to strain energy while quantifying uncertainty, achieving reliable predictions for unseen in-distribution data. However, its predictive accuracy and confidence decline for out-of-distribution microstructures with morphologically distinct attributes. Notably, while the model performs well in interpolating between small and medium pore sizes, it struggles with medium-to-large pore sizes due to significant differences in the deformation mechanisms of aerogels with larger pores.
Future work will focus on enhancing the framework by incorporating aerogel synthesis parameters as additional inputs to the WGAN-GP, enabling the discovery of novel morphologies that interpolate between training data for specific synthesis parameter combinations, as demonstrated in [37]. The BayesCNN will be further evaluated across a wider range of aerogel microstructures, investigating both interpolation and extrapolation capabilities. This includes assessing predictive accuracy for microstructures with diverse pore size distributions and integrating nonlinear finite element models based on continuum damage mechanics to capture complex deformation behaviors of ceramic aerogels. Additionally, the surrogate model will be refined through improved architecture design and hyperparameter selection using a hierarchical Bayesian inference framework [18]. This approach will identify the most plausible BayesCNN model under constraints of limited and uncertain training data, enhancing the framework reliable deployment in high-consequence material discovery scenarios.
This study tackles the computational challenges of model-driven materials discovery by developing a validated deep learning surrogate model for synthesis-microstructure-property relationships using limited and uncertain high-fidelity data. It also paves the way for advancing predictive accuracy and confidence in data-driven surrogate modeling for physical and engineering systems.

### v3

@@ 5 Conclusions
This study introduces an integrated deep learning surrogate modeling framework for efficient uncertainty propagation across the synthesis, microstructure, and properties of ceramic aerogel porous materials. The stochastic microstructure surrogate employs a scalable Wasserstein Generative Adversarial Network with Gradient Penalty (WGAN-GP) to generate ensembles of morphologically consistent microstructures over arbitrarily large domains, trained on smaller domain size samples from Lattice Boltzmann simulations of the foaming process. For property prediction, a Bayesian Convolutional Neural Network (BayesCNN) is developed to estimate the probability distribution of weight parameters from finite element training data, enables not only prediction of the strain energy but also quantification of the associated uncertainty, providing a measure of confidence in the surrogate model’s predictions. Numerical results demonstrate that the WGAN-GP effectively generates microstructures consistent with the morphology of training data (in terms of pore size distribution and two-point correlation) across larger domains. The BayesCNN demonstrates accurate mapping from microstructure to strain energy for in-distribution samples, with uncertainty estimates that reflect the level of confidence in prediction. Notably, both prediction error and uncertainty (i.e., variance) increase with average pore size, consistent with the broader distribution of mechanical responses in larger-pore microstructures. However, the model exhibits reduced predictive accuracy and confidence for microstructures with pore sizes interpolating between the training classes.
Future work will focus on enhancing the surrogate modeling framework by incorporating aerogel synthesis parameters as additional inputs to the WGAN-GP, enabling the discovery of novel morphologies that interpolate between training data for specific synthesis parameter combinations, as demonstrated in [41]. The predictive capability of the BayesCNN surrogate will be further evaluated on a broader set of 3D aerogel microstructures with diverse pore size distributions. Recent work has demonstrated the effectiveness of advanced CNN architectures for predicting material properties from complex 3D porous microstructures (e.g., [59, 60]). Extending our framework to such 3D settings will be a natural next step. In particular, we are interested in applying Bayesian training to these 3D models to not only assess predictive accuracy but also evaluate the associated confidence levels in model predictions. In parallel, we will extend the current framework to incorporate nonlinear and nonlocal finite element simulations with micromechanical damage models, enabling more accurate predictions of deformation behavior beyond the linear elastic regime. Such size-dependent models will account for the influence of solid wall thickness and failure mechanisms critical to the mechanical strength of ceramic aerogels. To improve BayesCNN reliability, architecture and hyperparameter selection will be guided by a hierarchical Bayesian inference framework [23], enabling identification of the most plausible model given limited training data, supporting more trustworthy surrogate modeling. An additional direction to improve the predictive capability of surrogate models is the incorporation of governing physical laws into the training process through physics-informed CNNs, e.g., [61, 62, 63], that has been shown to enhance extrapolation behavior and reduce data requirements. However, extending this approach within a Bayesian framework introduces new challenges. Specifically, it requires careful construction of a composite likelihood that balances data fidelity with residual physics violations under uncertainty. Hyperparameter tuning becomes particularly critical to ensure the model appropriately weighs physical constraints without compromising fit to available data, e.g., [64, 65]. The FSLBM simulations will also be extended to more faithfully represent the ceramic aerogel foaming process. This includes integrating a homogeneous nucleation model [66] to probabilistically determine initial bubble configurations based on synthesis parameters. The nucleation model incorporates two key parameters: surface tension and critical bubble radius. Surface tension will be modeled as a decreasing function of CTAB concentration, accounting for surfactant-induced stabilization at bubble interfaces. Higher CTAB concentrations enhance disjoining pressure, promoting bubble generation and suppressing coalescence [27]. In parallel, urea concentration, affecting bubble formation through thermal hydrolysis, will be used to inversely scale the critical bubble radius, as it governs the pressure differential driving nucleation. Finally, while the current microstructure-property surrogate relies on standard CNN architectures, future work will explore more advanced deep learning models [67, 68, 69, 70, 71] with superior generalization capabilities. Embedding these models within a Bayesian inference framework will be investigated to evaluate their effectiveness for uncertainty-aware surrogate modeling of complex porous materials such as ceramic aerogels.
This study addresses key computational challenges in developing reliable deep learning surrogate models for synthesis-microstructure-property relationships. The proposed framework, along with its extensions, establishes a foundation for uncertainty quantification in data-driven surrogate modeling, enabling broad applications in computational materials science, including uncertainty propagation, materials design, and optimization.
Data availability
The dataset and python code pertinent to this paper is available at https://github.com/pce-lab/BayesCNNSurrogate.